Based on the data below, calculate the Average Propensity to Consume at a disposable income of $500 Aggregate Disposable Income Consumption $ billions) $ billions) so $80 $100 $200 $300 $400 $500 $160 $220 $300 $380 $460 0.80 O$80 0.08 0.92 2.5 pts D Question 31 If disposable income increases from $450 to $470 bi propensity to save (MPS)? llion and savings increases from $15 to $20 billion, what is the marginal 0.25 0.02
Answer #1

The average propensity to consume(APC) is total consumption/ total income.

given the disposable income is $500, consumption is $460. Therefore, APC = 460/500= 0.92

31)Marginal propensity to save(MPS)= Change in saving/ change in income

change in saving=20-15 =$5, change in income= 470-450= $20

MPS = 5/20= 0.25